### v1

# Ranked from Within: Ranking Large Multimodal Models for Visual Question Answering Without Labels

Abstract
As large multimodal models (LMMs) are increasingly deployed across diverse applications, the need for adaptable, real-world model ranking has become paramount. Traditional evaluation methods are largely dataset-centric, relying on fixed, labeled datasets and supervised metrics, which are resource-intensive and may lack generalizability to novel scenarios, highlighting the importance of unsupervised ranking. In this work, we explore unsupervised model ranking for LMMs by leveraging their uncertainty signals, such as softmax probabilities. We evaluate state-of-the-art LMMs (e.g., LLaVA) across visual question answering benchmarks, analyzing how uncertainty-based metrics can reflect model performance. Our findings show that uncertainty scores derived from softmax distributions provide a robust, consistent basis for ranking models across varied tasks. This finding enables the ranking of LMMs on real-world, unlabeled data for visual question answering, providing a practical approach for selecting models across diverse domains without requiring manual annotation.
1 Introduction
Large Multimodal Models (LMMs), such as LLaVA [46] and InstructBLIP [11], have demonstrated remarkable capabilities in handling a wide array of complex multimodal tasks, proving highly adaptable across diverse real-world applications [44, 6, 11, 8, 5, 31]. From addressing scientific challenges [47, 22] and performing optical character recognition [55, 52, 53] to identifying spatial positions of objects [70, 77], LMMs are increasingly essential in practical settings. As LMMs evolve, the need for rigorous evaluation metrics that accurately capture their capabilities and limitations has become more urgent. Thus, numerous benchmarks have been developed [47, 22, 52, 67, 53, 77, 80], aiming to provide reliable rankings and guide users in selecting models best suited for specific deployment scenarios.
However, these benchmarks rely on carefully curated datasets, which can require substantial resources to develop and label. For many users who may not have access to these resources, assessing model performance can be challenging. Additionally, standard evaluations are often dataset-centric, depending on fixed, human-labeled metrics that may not capture the full range of model capabilities necessary for diverse applications. As LMM tasks continue to diversify and expand, evaluating them effectively has become increasingly complex, with new applications frequently needing additional data curation and specialized capabilities.
Addressing this challenge is essential for developing adaptable and scalable evaluation methods that support the safe and effective deployment of LMMs across diverse real-world environments. As shown in Fig. 1, selecting the most suitable LMM from a range of available options is challenging in the absence of annotations. To tackle this, our work proposes a proxy ranking metric that closely aligns with true performance, enabling effective model comparison in label-scarce settings. In image classification, several studies have aimed to predict a model’s performance on out-of-distribution (OOD) datasets using in-distribution (ID) metrics [68, 54, 4]. Known as accuracy-on-the-line (AoL), this approach demonstrates a strong linear correlation between probit-scaled ID accuracy and OOD accuracy across different classifiers. For ranking LMMs, we can leverage the AoL phenomenon [54] as a starting point. However, defining ID performance for LMMs is often challenging, as training datasets may be proprietary and training tasks may differ from test tasks. Therefore, we utilize performance on existing benchmarks as a practical basis for ranking models in new testing domains.
To explore this further, this paper begins by conducting a correlation study between performance on different benchmarks to examine the effectiveness of AoL for ranking models. With the observed instability and inconsistency of AoL for ranking LMMs, we then investigate the effectiveness of prediction uncertainty as the ranking indicator. LMMs generate answers in an open-form by producing token sequences, where each token is sampled from a probability distribution over the vocabulary. This enables assessing model uncertainty using logits on different token positions. To this end, we evaluate LMMs from different series. They have different training frameworks, such as LLaVA-V1.5 [44] and InstructBLIP [11], different visual encoders (e.g., CLIP [61] and SigLIP [81]) and language models (e.g., Vicuna [69] and LLaMA [71]). We investigate three perspectives to explore model uncertainty, which are softmax probability, self-consistency and labeled proxy set. We evaluate the ranking performance on widely-adopted multimodal benchmarks, which span in diverse domains, including reasoning scientific questions, recognizing optical characters and identifying objects’ spatial positions. Our main observations are summarized as below:
We demonstrate that model performance on one dataset may not accurately reflect the model rankings on a different dataset (Section 4).
We evaluate various ranking methods and observe that their effectiveness is influenced by task characteristics (i.e., closed-form or free-form generation). Among the methods considered, NLLmin stands out as the most effective (Section 5).
When examining correlations in model performance across different dataset pairs, we observe that variations in text prompt similarity align more closely with performance correlation changes than variations in image feature similarity (Section 6).
2 Related Work
Unsupervised Model Ranking.
The goal of this task is to rank and select a best performant models without the access to the data annotations of target environments. The research on this task can date back to Forster et al. [14], and was further investigated in various tasks: (1) outlier detection [84, 83]; (2) image classification [36, 73, 85, 4, 54, 64]; (3) time series anomaly detection [78]; (4) multivariate anomaly detection in manufacturing systems [13], etc.
We discuss the most relevant two lines of research as follows. Miller et al. [54] introduce an accuracy-on-the-line (AoL) phenomenon where strong linear correlation between probit-scaled in-distribution (ID) accuracy and out-of-distribution (OOD) accuracy across a variety of ML models. This means ID accuracy serves as a good indicator of model performance for a target domain. Shi et al. [64] revisit the established concept of lowest common ancestor (LCA) distance, which measures the hierarchical distance between labels and predictions within a predefined class hierarchy. By the observed linear correlation between ID LCA distance and OOD accuracy, it is viable to select a model based LCA distance of predictions. This work differs from prior study that we focus on ranking LMMs. This group of models makes predictions in a significantly different way from conventional classification models, which also results in different evaluation protocols.
Uncertainty estimation for LLMs and LMMs.
Uncertainty estimation aims to measure an ML model’s confidence level of a certain prediction [20]. There are two major types of uncertainty one can model: aleatoric and epistemic uncertainty [33]. Aleatoric uncertainty refers to the noise inherent in the observations; while epistemic uncertainty accounts for uncertainty in the model parameters. To quantify uncertainty, uncertainty estimation are roughly categorized into four groups: (1) single deterministic method [20, 21]; (2) model ensembling [39]; (3) Bayesian method [16, 10, 38]; (4) test-time augmentation [48, 75]. Single deterministic methods calculate prediction uncertainty based on one forward pass within a deterministic ML model. By contrast, Bayesian methods only leverage the inherent stochasticity of an ML model (e.g.,, dropout layer in deep neural networks [16]). We primarily use (1) and (3) for ranking LMMs due to their simplicity in implementation and model- and domain-agnostic.
Evaluation & Benchmarking LMMs.
The rapid development of LMMs has greatly propelled advancements in multimodal models, showcasing significant improvements in their perception and reasoning abilities. This shift has rendered traditional benchmarks, which focus solely on isolated task performance [32, 3], inadequate for thorough evaluation. Researchers have introduced new benchmarks aimed at evaluating LMMs across a broad spectrum of multimodal tasks [19, 43, 63]. However, many benchmarks lack detailed assessments and robust evaluation measures. Recent studies [80, 77] highlight the need for more comprehensive benchmarks to evaluate the reasoning and understanding abilities of LMMs effectively.
3 Task Formulation
Task Definition.
Let a multi-modal task be represented by a dataset , where and denote the input prompt and the corresponding answer for the -th sample, respectively. We have access to large multimodal models (LMMs), denoted as . Each LMM , with pre-trained weights , generates a sequence of tokens from the input prompt via a decoding process: , where represents the -th generated token. To assess the performance of each LMM, this task employs an evaluation metric (e.g., accuracy) that determines the ground-truth performance by comparing the generated sequences with the ground-truth answers .
The objective of this paper is to develop methods for computing scores for each LMM without requiring access to task-specific data annotations. Ideally, these computed scores should closely correlate with ground-truth performance, allowing us to rank and select LMMs based on their performance using only these scores.
Evaluation Metric.
We use Spearman’s rank correlation coefficient () [34] to assess the monotonic relationship between the computed scores and model performance. Additionally, we calculate the weighted version of Kendall’s rank correlation () [66], which has been shown to effectively identify top-ranked items of interest [79]. Both coefficients range from , where values close to or indicate strong negative or positive correlations, respectively, and a value of denotes no correlation.
4 Uncertainty for Ranking LMMs
In this section, we discuss the unique characteristics of ranking various Large Multimodal Models (LMMs) compared to conventional ML models. We then present three distinct approaches that leverage uncertainty in model predictions for ranking.
4.1 What makes ranking LMMs interesting?
Unique Challenges and Potential Methods for Ranking LMMs.
While LMMs can be considered a subset of machine learning (ML) models, ranking them introduces unique challenges not present in traditional ML models. Below, we discuss the unique characteristics of LMMs and the challenges they bring to the model-ranking process.
Scale and Architectural Complexity. LMMs, often pre-trained with billions of parameters (e.g., LLaVA-V1.5-7B [46] with billion parameters), present significant challenges for traditional “white-box” analyses. Techniques such as examining neuron activations [2], analyzing attention patterns [17, 74], or assessing weight norms across layers [57] are computationally and manually prohibitive due to their sheer scale.
Limited Transparency. LMMs also exhibit transparency challenges. First, these models are typically trained on large datasets of instruction fine-tuning samples, which may be publicly accessible or proprietary data. As a result, risk assessment methods that require access to training data (e.g., uncertainty estimation [20]) are either unsuitable or should be adjusted. Additionally, although some LMMs may disclose their final model weights, other information, such as training loss and intermediate checkpoints, often remains undisclosed. The lack of the access to training data and training details limits the use of techniques in the unsupervised accuracy estimation [12, 72] and scaling law [28, 29].
Using Accuracy-on-the-Line (AoL)?
Given these challenges, one may consider leveraging the AoL phenomenon as a potential method for ranking LMMs. Miller et al. [54] describe AoL as a strong linear correlation between probit-scaled in-distribution (ID) accuracy and out-of-distribution (OOD) accuracy across various ML models. This suggests that ID accuracy could serve as a reliable predictor of OOD performance, making AoL an intuitive starting point for selecting LMMs in target testing environments.
However, we suggest that AoL may not be suitable in the context of ranking LMMs for real-world deployment. First, obtaining ID performance data is often impractical, as LMMs are typically trained on large, sometimes proprietary datasets, limiting direct access to ID metrics. Second, while performance on existing benchmarks is more readily available, using this data to rank models for new deployment environments poses significant challenges. Figure 2 illustrates the correlation between proxy benchmark performance and target testing environments, revealing substantial variability in correlation strength, ranging from as high as to a negative correlation of . This inconsistency highlights the unreliability of using benchmark performance as a sole ranking criterion. Third, relying solely on proxy benchmark performance fails to capture the unique statistical characteristics of target testing datasets. AoL tends to rank models identically across different environments, regardless of the actual deployment context, making it less suitable for selecting LMMs intended for diverse, fine-tuning-free environments.
Last, a practical alternative could involve annotating a subset of the target environment and using the performance on this labeled subset as a ranking indicator, if the budget allows. This method might be useful in scenarios with non-stationary distributions, as periodic re-annotation may be needed to reflect changes in the testing environment. However, detecting such shifts requires carefully designed methods, and in specialized domains like medical image analysis [62] or harmful content detection [35], annotation can be both time-consuming and require expert knowledge.
What Are the Possible Methods?
As model size and architectural complexity increase, traditional methods become less feasible due to the substantial computational demands and the need for more sophisticated algorithm design. Techniques that leverage scaling laws [30] also require multiple models from the same family, but with varying sizes, to train a regressor that estimates model performance. In model ranking contexts, computational overhead grows substantially with the number of models, making such methods impractical for large-scale deployment.
In contrast, model outputs—such as the logits used for token prediction and the generated tokens—are readily accessible without requiring deep examination of the model architecture or complex information extraction techniques. Therefore, we focus primarily on methods that utilize the model’s prediction logits and output strings to rank LMMs.
4.2 Uncertainty by Softmax Probability
Unlike models used in image classification, where predictions correspond to the entry with the highest softmax probability, LMMs generate answers in an open-form manner by producing sequences of tokens. However, each token generation can still be viewed as a classification process, where the model selects the token with the highest probability over its entire vocabulary. This allows us to assess model uncertainty based on the logits at each token position.
For token-level uncertainty, we can focus on two specific positions: the first token and the penultimate token (i.e., the token preceding the end-of-sequence token), as illustrated in Figure 3. The logit of the first token reflects the model’s initial response to the input prompt, while the logit of the penultimate token captures the model’s understanding of both the prompt and the generated response. The uncertainty associated with these two positions may provide insight into the model’s overall confidence in answering the question. Beyond token-level uncertainty, sentence-level uncertainty can be calculated by aggregating uncertainty across all tokens in the sequence, as suggested in [50, 23]. Specifically, sentence-level uncertainty can be quantified using the mean or the maximal negative log likelihood values across the entire generated sequence:
where is the likelihood of the word generated by the LMM at the -th token of the -th sentence and is the number of tokens generated for this sentence. Equation 1 quantifies a sentence’s uncertainty by identifying its least likely token, while the average per-token uncertainty (Equation 2) normalizes the log probability across sequences of varying lengths, allowing for length-independent comparisons of uncertainty [49, 56]. In addition, Equation 2 is closely related to [25, 51], which provides a comparable measure of model quality.
Alternatively, entropy can also be used instead of negative log likelihood to assess uncertainty. In the context of model ranking, we adopt normalized entropy to account for varying vocabulary sizes across models, thus scaling entropy to the interval and making it comparable across different models. The normalized entropy is given by:
where is the likelihood of the word being generated at the -th token of the -th sentence, and is the set of all possible words in the vocabulary.
There are eight variants of output probability-based methods, denoted as NLLF, NLLP, NLLmin, NLLavg, EntF, EntP, Entmax, and Entavg. “NLL” and “Ent” represent negative log likelihood and entropy, respectively, while “F” and “P” refer to the first and penultimate tokens.
4.3 Uncertainty by Self-Consistency
Another approach involves examining the non-deterministic generations produced by models. The core intuition is that a more accurate model will produce predictions closely aligned with the original answer, while a less accurate model may yield more varied responses with each inference. In the case of LMMs, the temperature parameter controls the randomness of predictions: a temperature of zero forces deterministic predictions, where the model selects only the token with the highest probability. When is larger than , the model generates tokens more stochastically, selecting tokens based on probabilities above a threshold. As increases, the model becomes increasingly stochastic in its token generation [7, 9, 23].
To analyze the consistency in these stochastic predictions, we explore two common methods: BLEU [59] and BERTScore [82]. BLEU is a n-gram-based metric that evaluates the similarity of generated sequences to the reference answer, while BERTScore uses a pre-trained language model to embed answers and measures similarity in embedding space. Then, we use the mean value of similarities to represent the consistency for the input sample, which can be denoted as , where is the number of stochastic inferences, is the similarity function (e.g., BLEU), and and are the -th stochastic prediction and the original answer, respectively. Following the practice outlined in [7, 9, 23], we collect five stochastic inferences per sample and set to maintain a relatively high degree of stochasticity in LMM generation while keeping compute overhead manageable. We adopt a unigram BLEU and denote the methods using BLEU and BERTScore as SampleBLEU and SampleBERT, respectively. Furthermore, BERTScore tends to assign high similarity scores between single letters, such as “A” and “B”, which should be considered very different in the contexts of MCVQ. For example, BERTScore gives a high similarity score of for “A” and “B”, making it ineffective for distinguishing model performance. To address this, we modified the response to include the full answer text (e.g., “A. North America”), denoted as Sample*BERT.
4.4 Uncertainty With a Labeled Proxy Dataset
Beyond the uncertainty score calculated directly on the target dataset, we also explore scores obtained using a proxy labeled dataset. In [18], average thresholded confidence (ATC) has been proposed. ATC calculates a threshold on a validation set (e.g., CIFAR-10 [37] validation set) and considers an image correctly classified on a new dataset with the same task if its confidence score exceeds the threshold. With the derived threshold, model performance on the target domain is estimated by the proportion of samples with confidence scores higher than the threshold. For LMMs, we use an existing benchmark as a proxy set to calculate . The calculation of ATC is formulated as follows:
where is a binary indicator function, represents an uncertainty estimation method, which we use NLLmin here. ATC provides both a ranking of model performance and an estimate of the expected performance.
Method | MCVQ | VQA | ||||||||||||||||
SQA-I | AI2D | MMMU | RWQA | ChartQA | OCRVQA | TextVQA | DocVQA | Average | ||||||||||
AoL | ||||||||||||||||||
NLLF | ||||||||||||||||||
NLLP | ||||||||||||||||||
NLLmin | ||||||||||||||||||
NLLavg | ||||||||||||||||||
EntF | ||||||||||||||||||
EntP | ||||||||||||||||||
Entmax | ||||||||||||||||||
Entavg | ||||||||||||||||||
SampleBLEU | ||||||||||||||||||
SampleBERT | ||||||||||||||||||
Sample*BERT | ||||||||||||||||||
ATC |
5 Experiments
5.1 Experiment setup
Benchmarks.
We consider both multiple choice visual question (MCVQ) and visual question answering (VQA) benchmarks. MCQ and VQA are both types of question-answering formats for evaluating LMMs. The difference lies on that the former is provided with several answer options, out of which only one (or sometimes more) is correct, the model is only required to select an answer option(s). In contrast, the latter is usually open-ended, models may generate answers freely. We consider widely-adopted MCVQ and VQA benchmarks. They are: (1) the subset of ScienceQA [47] with images (SQA-I) and AI2D [22] which assess LMMs’ scientific knowledge; (2) ChartQA [52], OCRVQA [55], TextVQA [67] and DocVQA [53] to examine their ability to recognize optical character; (3) RealWorldQA (RWQA) [77] which evaluates LMMs’ vision-centric capability; (4) MMMU [80] which assays LMMs on multi-discipline tasks that demand college-level subject knowledge. Note that, SQA-I, AI2D, RWQA and MMMU are MCVQ datasets, while the others are VQA.
Models.
We consider two different settings. For the first setting, the goal is to choose the best LMM over all different series of LMMs. We include LLaVA-V1.5 [44], ShareGPT4V [6], LLaVA-NeXT [45], InstructBLIP [11], LLaVA-NeXT-Interleave [42], LLaVA-OneVision [40], Eagle [65], mPLUG-Owl [41], InternVL [8], PaliGemma [5], Mantis [27]. In total, we have collected different models, which all can be accessed on Huggingface [76].
5.2 Main Observations
Accuracy-on-the-line could be a less reliable indicator for ranking LMMs in new domains. Table 1 summarizes the ranking performance of all methods across eight tasks. The AoL performance is calculated as the average correlation strength when using the other seven domains to predict model rankings in the target domain. We observe that AoL does not achieve consistently high correlation with model performance in seven out of the eight tasks. Although AoL shows strong results on RealWorldQA, where it performs best, uncertainty-based methods also demonstrate high correlation strength. For instance, NLLmin exhibits a Spearman’s lower but a weighted Kendall’s higher than AoL. These findings suggest that relying on existing benchmarks alone to select models for safe deployment can be risky, as they may not capture the statistical characteristics of target domains.
The choice of token matters for output-probability-based methods’ ranking capabilities.
In generated sequences, four ways of token usages can be used to extract model uncertainty. For the two on specific positions, i.e., the first and penultimate tokens, their predictive performance depends on the task type (i.e., MCVQ vs. VQA). The uncertainty associated with generating the first token is more indicative for MCVQ tasks, while the penultimate token generally proves more predictive for VQA tasks. This difference is due to the nature of MCVQ tasks, which often prompt models to generate only a single option letter (e.g., “A”), making the first token crucial for evaluating response accuracy. In contrast, VQA tasks require open-form responses consisting of multiple tokens, making the penultimate token-reflecting the model’s understanding of both the question and the answer-more informative.
For the two span within the generated output, i.e., NLLmin and NLLavg, we find they are more stable and less task-specific compared to individual tokens. While NLLmin and Entmax (reflecting the least confident token) is more effective for MCVQ tasks, both methods perform similarly on VQA tasks. The superior performance of using the first token and the least confident token as ranking indicators for MCVQ suggests that a single token (typically the option letter) holds significant meaning. Considering all potential tokens can sometimes result in an unexpectedly higher confidence level, as LMMs may generate the complete answer alongside the option letter (e.g., “B. Columbia”). The tokens following the option letter often have high confidence levels, leading to an overall increase in estimated uncertainty. This tendency results in models generating entire answers with lower uncertainty scores, yielding higher ranks. These findings underscore the importance of further exploration into the optimal token positions for uncertainty estimation in LMMs. One potential approach is leveraging a language model to identify the position of the option letter and use its softmax probability as the uncertainty measure for the complete response.
Negative log likelihood (NLL) is more stable and predictive for model ranking than normalized entropy.
For seven out of eight tasks, NLL consistently shows stronger correlation with model performance than entropy. Both NLL and entropy leverage softmax probabilities during token generation; however, NLL focuses on the maximal probability, while entropy considers all entries. The lower correlation strength of normalized entropy may stem from the significantly different vocabulary sizes of various LMMs. For instance, LLaVA-V1.5-7B has a vocabulary of k tokens, while PaliGemma-3B-mix has a much larger vocabulary of k tokens. Despite being normalized by vocabulary size, entropy may be still susceptible to noise than NLL. Future work could include applying dimension reduction techniques or limiting consideration to the top- most probable tokens or a pre-defined set of tokens.
Sample-based methods are strong candidates for model ranking without intrinsic access to LMMs.
The performance of sample-based methods is influenced by the nature of the task. They yield higher correlation scores on VQA tasks compared to MCVQ tasks. In VQA, models typically produce more varied responses, making BLEU and BERTScore effective for capturing uncertainty. However, MCVQ tasks constrain models to select from a pre-defined set of answers. While non-zero temperature introduces some randomness, the variation between stochastic inferences is limited. Additionally, we also find that Sample*BERT has higher correlation scores than SampleBERT on four MCVQ benchmarks and suggests that developing more advanced algorithms to capture uncertainty from multiple stochastic inferences could be beneficial.
Although sample-based methods generally show lower overall performance compared to other methods, they remain competitive without the need to examine model architecture or intrinsic states. This highlights their potential for ranking LMMs that are accessed through APIs, such as GPT-4V [1]. Additionally, the performance of sample-based methods can be enhanced by tuning the temperature parameter and increasing the number of stochastic inferences, which could yield more reliable results.
ATC indicates model ranking to some extent, but the correlation strength is influenced by the choice of proxy dataset.
We compute ATC performance as the average correlation when using the other seven datasets as proxy datasets. Our analysis reveals the choice of proxy dataset is critical, as the scale of uncertainty calculated on different datasets can vary. This variation can lead to an inaccurately estimated threshold for determining instance correctness. A potential improvement involves adopting uncertainty calibration methods, such as temperature scaling [20], to calibrate model uncertainty onto a consistent scale.
6 Discussion
Method | MCVQ | VQA | Average | ||
AI2D | MMMU | TextVQA | ChartQA | ||
AoL | |||||
NLLmin | |||||
NLLavg | |||||
Entmax | |||||
Entavg | |||||
SampleBLEU | |||||
ATC |
Ranking LMMs from the same series.
So far, all experiments have focused on selecting LMMs from different model series. However, in some scenarios, the objective is to choose a training recipe that yields better-performing LMMs within the same model series. These models may be trained with additional fine-tuning steps, varied data augmentations, different training sources, or different language models (e.g., Vicuna and Mistral [26]) and visual encoders, such as CLIP and DINOv2 [58]. For our analysis, we use the LLaVA series due to its widespread adoption. Specifically, we employ 15 different LLaVA prismatic (LLaVA-pri) models [31]. We report results on four datasets in Table 2: AI2D, MMMU, TextVQA, and ChartQA, with AI2D and TextVQA being used to evaluate the performance of different LLaVA variants in the original paper. The AoL and ATC performance metrics are computed in the same manner as described in Section 5, using the average correlation across other three datasets.
Our findings align with those from ranking different series of LMMs. First, using model performance on a single existing dataset may not accurately reflect model rankings on a different domain, even when the models are trained with a similar pipeline and minor variations. Second, we observe that ranking different LLaVA-pri models in MCVQ presents a significant challenge, as the variance in model performance on MCVQ tasks is lower compared to VQA tasks. For instance, the performance gap on AI2D is only , whereas the gap on TextVQA exceeds . This indicates that ranking methods must capture subtle differences between models. We find that SampleBLEU remains effective, while NLL and Entropy may not capture these nuances accurately. Additionally, the decrease in correlation between uncertainty estimated by softmax probability suggests that although modifications to the training pipeline may have a slight effect on performance, they can lead to significant variations in the confidence levels for generation. This finding underscores the importance of assessing LMMs more comprehensively, beyond accuracy alone.
Analysis on the weak correlation of AoL.
Figure 2(b) illustrates the correlation of model performance across different benchmarks. We observe that while TextVQA, ChartQA, DocVQA, and OCRVQA all aim to assess the capability of LMMs to recognize optical characters, the correlation between them varies significantly. Specifically, model performance on TextVQA, ChartQA, and DocVQA shows strong correlations, whereas performance on OCRVQA consistently exhibits low correlation with the other three datasets. To explore whether the images or texts within these datasets influence the correlation strength of model performance, we utilize CLIP-L-14-336 [61] to extract image and text embeddings. We then use the Fréchet distance (FD) [15] to measure the similarity between datasets based on these features.
Figure 4 presents a correlation study between the FD of dataset pairs and the correlation strength of model performance on those datasets. We observe a strong correlation for FDs computed using text input features, while FDs measured by image features show a weaker correlation. This suggests that text input similarity is likely a more influential factor for model performance correlation than image similarity. Additionally, orange stars are used to label the points representing the lowest FD for each dataset. Moreover, OCRVQA shows a low correlation strength with other datasets (Fig. 2). The closest dataset in prompt feature space is TextVQA, with a FD of 124, which is considerably higher than the lowest FDs observed between other dataset pairs, typically around 70 or lower. This difference sheds light on the weak model performance correlation between OCRVQA and other OCR-focused datasets.
Method | MCVQ | VQA | ||
AI2D | MMMU | TextVQA | ChartQA | |
#Samples | ||||
samples | ||||
NLLmin |
Ranking LMMs by a labeled subset of target domain.
Table 3 presents the correlation strength between model performance on labeled instances in the target domain and the overall performance across the entire dataset. We observe that a small labeled set can provides good rankings. Yet, while the sampled data are not representative, such way may not reliably capture the model ranking for the full dataset [60] In contrast, the method NLLmin gives consistently high correlation, highlighting the potential of uncertainty-based methods for effective model ranking.
7 Conclusion
This work studies an interesting question: ranking large multimodal models without the access to target domain labels. Our analysis begins by identifying weak correlations in model performance across different domains, leading us to explore uncertainty scores estimated from model predictions. We evaluate 45 LMMs on visual question answering, testing various training frameworks, visual encoders, and language models. Our experiments reveal that negative log likelihood serves as a useful performance indicator, and while stochastic sampling can be helpful, it is less effective for multiple-choice tasks, where it requires repeated inferences and careful temperature tuning. This study establishes a baseline for using uncertainty in model ranking, paving the way for further research
Several exciting directions remain open. Beyond uncertainty scores, test-time augmentation and semantic entropy may offer additional insights. Our findings suggest that text feature dissimilarity has a stronger impact on weak performance correlations across datasets than image feature dissimilarity. Exploring other factors to design more effective multimodal benchmarks presents an intriguing avenue.